### v2

# Measuring and Analyzing Intelligence via Contextual Uncertainty in Large Language Models using Information-Theoretic Metrics

Abstract
Large Language Models (LLMs) excel on many task-specific benchmarks, yet the mechanisms that drive this success remain poorly understood. We move from asking what these systems can do to asking how they process information. Our contribution is a task-agnostic method that builds a quantitative Cognitive Profile for any model. The profile is built around the Entropy Decay Curve—a plot of a model’s normalised predictive uncertainty as context length grows. Across several state-of-the-art LLMs and diverse texts, the curves expose distinctive, stable profiles that depend on both model scale and text complexity. We also propose the Information Gain Span (IGS) as a single index that summarises the desirability of a decay pattern. Together, these tools offer a principled way to analyse and compare the internal dynamics of modern AI systems.
1 Introduction
Intelligence defies a single, agreed-upon definition Neisser (1979); Sternberg (2004). Yet recent LLMs display behaviours that look intelligent, reaching human-level scores on complex tasks. If the concept itself is elusive Bommasani et al. (2021); Bender et al. (2021), how can we understand systems that seem to embody it?
LLMs give us a unique opening. For any context, they output a full probability distribution over the next token Brown et al. (2020); Bishop and Nasrabadi (2006). These distributions, derived from the model’s internal logits, let us study information processing directly.
We ground our analysis in Shannon’s information theory Shannon (1948). Prior work has used information measures to trace information flow in deep nets Tishby et al. (2000); Shwartz-Ziv and Tishby (2017) and to gauge LLM confidence Kadavath et al. (2022); Jiang et al. (2020). We build on this by noting that intelligent behaviour balances two extremes: rigid determinism, which lacks creativity, and pure randomness, which is noise. A capable system should lock in on clear evidence yet stay flexible when context is thin. We call this capacity Adaptive Predictive Modulation.
To measure it, we compute two entropy-based quantities for a fixed context length $k$:
  • Average conditional entropy $h_{k}$: residual uncertainty after seeing $k$ tokens.
  • Entropy of the average distribution $H_{k}$: diversity of possible outputs averaged over all contexts of length $k$.
(Definitions follow in Section 2). From these we form the Length-Conditional Uncertainty Index
$u_{k}:=\frac{h_{k}}{H_{k}},\hskip 17.00024pt0\leq u_{k}\leq 1,$
which places the model on a determinism–randomness scale for each $k$.
By tracing $u_{k}$ against $k$, we obtain the Entropy Decay Curve—a quantitative Cognitive Profile [^1] Perplexity (PPL) relates to conditional entropy via $h=\log_{2}(\text{PPL})$, but PPL is usually reported as a single corpus-level score. The index $u_{k}$ is new: it also uses $H_{k}$, which PPL ignores.. The curve’s starting point, slope, and floor capture a model’s information-handling strategy. We present the first empirical study of these profiles for leading LLMs, showing how they reveal both genuine capability and issues such as data contamination Carlini et al. (2022); Achiam et al. (2023).
The rest of the paper defines the metrics, describes the experimental design, and analyses the resulting profiles.
2 Definitions and Metrics
We begin by formalising the probability distributions produced by a Large Language Model (LLM) and the information-theoretic measures derived from them. These definitions underlie all subsequent analyses of predictive uncertainty.
2.1 Predictive Distribution of an LLM
Consider an auto-regressive LLM that, given a context of $k$ tokens, outputs a probability distribution over its full vocabulary. Let
$X=(t_{1},t_{2},\dots,t_{k})$
be such a context and let $\mathcal{Y}$ denote the vocabulary (typically $|\mathcal{Y}|\approx 128\,256$). The model returns the conditional distribution
$p(Y\mid X),$
where, for any token $y\in\mathcal{Y}$,
$p(y\mid X)=\Pr\bigl(\text{next token}=y\mid X\bigr).$
Crucially, $p(Y\mid X)$ refers to the entire probability vector, not just a sampled outcome. The vector is obtained by applying the softmax function to the model’s logits, the standard way to convert raw scores into a valid probability distribution in multi-class generation tasks Goodfellow et al. (2016); Vaswani et al. (2017).
2.2 Empirical Estimation of Entropies
We estimate conditional and marginal entropies for a fixed context length $k$ by sampling $N$ context windows $\{x_{1},x_{2},\dots,x_{N}\}$ from a representative corpus.
2.2.1 Conditional Entropy $\boldsymbol{(h_{k})}$
For each context $x_{i}$ we compute its conditional entropy
$h_{x_{i}}$ $:=H(Y\mid X=x_{i})$
$=-\sum_{y\in\mathcal{Y}}p(y\mid X=x_{i})\,\log_{2}p(y\mid X=x_{i}).$
The average conditional entropy for length $k$ is the sample mean
$h_{k}:=\frac{1}{N}\sum_{i=1}^{N}h_{x_{i}},$
which represents the model’s mean predictive uncertainty after reading $k$ tokens.
2.2.2 Marginal Entropy $\boldsymbol{(H_{k})}$
Exact marginal entropy requires averaging over all contexts and is intractable Cover (1999). Instead, we form the average predictive distribution
$\bar{p}_{k}(y):=\frac{1}{N}\sum_{i=1}^{N}p(y\mid X=x_{i}),\hskip 17.00024pty\in\mathcal{Y},$
and define the empirical marginal entropy
$H_{k}:=-\sum_{y\in\mathcal{Y}}\bar{p}_{k}(y)\,\log_{2}\bar{p}_{k}(y).$
This value measures the overall diversity of tokens the model finds plausible, given $k$-token contexts, and serves as an upper bound for $h_{k}$.
2.3 The Cognitive Profile: Uncertainty Index and Entropy Decay Curve
Length-Conditional Uncertainty Index
For each $k$ we form
$u_{k}:=\frac{h_{k}}{H_{k}},\hskip 17.00024pt0\leq u_{k}\leq 1,$ (1)
called the Length-Conditional Uncertainty Index. It rescales residual uncertainty by the model’s potential output diversity at that context length.
Entropy Decay Curve (EDC)
Plotting $u_{k}$ against $k$ (e.g., $k\in\{3,9,30,\dots\}$) yields the Entropy Decay Curve. The curve’s starting height, rate of decline, and eventual plateau together provide a quantitative Cognitive Profile of the model’s information-processing behaviour.
3 Experimental Setup
This section describes the models, corpora, and procedures used to estimate the Entropy Decay Curves (EDCs).
3.1 Models
To compare information processing across architectures and sizes, we test three public LLMs. All are quantised to GGUF format with Q4_K_M weights for consistent, memory-efficient inference:
  • Llama 3.3 70.6B — 70.6-billion parameters;
  • DeepSeek-R1 8.19B — 8.19-billion parameters in the Qwen3 line Yang et al. (2025);
  • Qwen 2.5 7.62B — 7.62-billion parameters in the Qwen2 family Yang et al. (2024).
3.2 Corpora
All evaluations use three public-domain texts from Project Gutenberg [^2] https://www.gutenberg.org/ebooks/11333https://www.gutenberg.org/ebooks/4300444https://www.gutenberg.org/ebooks/48433 :
  1. 1.
    Alice’s Adventures in Wonderland — analysis begins at “CHAPTER I. Down the Rabbit-Hole…”.
  2. 2.
    Ulysses — analysis begins at “– I – [1] Stately, plump Buck Mulligan…”.
  3. 3.
    Kant’s Critique of Judgement — analysis begins at “PREFACE We may call the faculty of cognition from principles…”.
Project Gutenberg headers, licences, and tables of contents are removed; the main text is left unchanged.
3.3 Implementation Details and Procedure
Software stack.
  • llama-cpp-python for quantised inference
  • NumPy for array operations
  • scipy.special for numerically stable softmax and entropy
Sliding-window evaluation.
  • Window lengths $k\in\{3,9,30,90,300,600\}$
  • Samples per $k$ $N=1000$
  • Tokens needed $1600=\max(k)+N$
  • Model context n_ctx =2048=2048 with full GPU off-load (n_gpu_layers =1=-1)
Algorithm.
For each model and each window size $k$:
  1. 1.
    Tokenise the first 1600 tokens of the chosen corpus.
  2. 2.
    For $i=1\dots N$:
    1. (a)
      Reset the model state.
    2. (b)
      Extract context $x_{i}=(t_{i},\dots,t_{i+k-1})$.
    3. (c)
      Run the model on $x_{i}$ to obtain logits.
    4. (d)
      Compute the conditional entropy $h_{x_{i}}$; store $p(Y\mid X=x_{i})$.
  3. 3.
    After $N$ iterations compute
    • $h_{k}$ — mean conditional entropy,
    • $H_{k}$ — entropy of the averaged distribution,
    • $u_{k}=h_{k}/H_{k}$ — Length-Conditional Uncertainty Index.
The full pipeline is repeated for every model and every $k$, producing the data required to plot each Cognitive Profile.
4 Results
We now quantify the predictive uncertainty exhibited by the three language models across multiple context lengths and corpora. The analysis centres on the conditional entropy $h_{k}$, the marginal entropy $H_{k}$, and their ratio $u_{k}$.
4.1 Empirical Data and Initial Observations
Tables 1, 2, and 3 report $h_{k}$, $H_{k}$, and the derived uncertainty index $u_{k}$ for all models on Ulysses, Alice’s Adventures in Wonderland, and Kant’s Critique of Judgement, respectively. The primary metric $u_{k}$ expresses residual prediction uncertainty normalised by the model’s potential output diversity at each context length $k$.
Table 1: Conditional entropy $h_{k}$, average‐distribution entropy $H_{k}$ (bits), and their ratio $h_{k}/H_{k}$ for each model on the Ulysses corpus.
Model Metric k=3 9 30 90 300 600
Llama 3.3 70.6B $h_{k}$ 11.2112 5.4613 1.9958 0.5029 0.2583 0.2160
$H_{k}$ 13.5514 10.5765 9.1482 8.3146 8.1838 8.1527
$h_{k}/H_{k}$ 0.8273 0.5164 0.2182 0.0605 0.0316 0.0265
DeepSeek-R1 8.19B $h_{k}$ 5.7411 4.6142 3.9565 3.7438 3.6196 3.5586
$H_{k}$ 10.2484 9.9338 9.6758 9.4896 9.3911 9.3497
$h_{k}/H_{k}$ 0.5602 0.4645 0.4089 0.3945 0.3854 0.3806
Qwen2.5 7.62B $h_{k}$ 6.3178 5.0146 4.3945 4.2200 3.8902 3.6960
$H_{k}$ 10.7054 10.0853 9.7528 9.5351 9.3776 9.2664
$h_{k}/H_{k}$ 0.5902 0.4972 0.4506 0.4426 0.4148 0.3989
Table 2: Conditional entropy $h_{k}$, average‐distribution entropy $H_{k}$ (bits), and their ratio $h_{k}/H_{k}$ for each model on the Alice’s Adventures in Wonderland corpus.
Model Metric k=3 9 30 90 300 600
Llama 3.3 70.6B $h_{k}$ 11.6270 5.0797 1.4042 0.1720 0.1086 0.1161
$H_{k}$ 13.4119 10.0430 8.4512 7.8135 7.7283 7.7352
$h_{k}/H_{k}$ 0.8669 0.5058 0.1662 0.0220 0.0141 0.0150
DeepSeek-R1 8.19B $h_{k}$ 5.4991 4.3606 2.6554 1.3696 1.2106 1.3834
$H_{k}$ 9.7100 9.3782 8.8974 8.3018 8.1619 8.1946
$h_{k}/H_{k}$ 0.5663 0.4650 0.2984 0.1650 0.1483 0.1688
Qwen2.5 7.62B $h_{k}$ 6.3029 4.7156 2.4262 0.7460 0.4949 0.5389
$H_{k}$ 10.2246 9.6059 8.7706 8.0662 7.9394 7.9347
$h_{k}/H_{k}$ 0.6164 0.4909 0.2766 0.0925 0.0623 0.0679
Table 3: Conditional entropy $h_{k}$, average‐distribution entropy $H_{k}$ (bits), and their ratio $h_{k}/H_{k}$ for each model on the Kant’s Critique of Judgement corpus.
Model Metric k=3 9 30 90 300 600
Llama 3.3 70.6B $h_{k}$ 11.5800 5.8695 3.1620 1.6728 1.4296 1.5319
$H_{k}$ 13.4347 10.2062 8.7336 8.0251 7.8971 7.8724
$h_{k}/H_{k}$ 0.8619 0.5751 0.3621 0.2084 0.1810 0.1946
DeepSeek-R1 8.19B $h_{k}$ 5.6830 4.7242 3.6155 2.9959 2.8928 2.9508
$H_{k}$ 9.8826 9.4601 8.8292 8.4820 8.3635 8.3338
$h_{k}/H_{k}$ 0.5751 0.4994 0.4095 0.3532 0.3459 0.3541
Qwen2.5 7.62B $h_{k}$ 6.4208 5.2071 3.7593 3.2637 3.2129 3.2450
$H_{k}$ 10.3692 9.6701 8.7061 8.4114 8.3495 8.3070
$h_{k}/H_{k}$ 0.6192 0.5385 0.4318 0.3880 0.3848 0.3906
Monotonic Decline in Uncertainty
For every model and corpus, the uncertainty index $u_{k}$ decreases as the context length $k$ grows. Models are most uncertain at $k=3$; confidence rises at $k=9$ and continues to improve through $k=30$ and beyond. This uniform trend confirms that additional context restricts the set of plausible next tokens.
Diminishing Returns from Longer Context
Although $u_{k}$ falls steadily, its decline slows once the context becomes long. The drop from $k=3$ to $k=30$ is much larger than the drop from $k=300$ to $k=600$. Thus most predictive benefit comes from early context, while very long windows add only modest gains.
4.2 Model-Specific Cognitive Profiles
Figures 1, 2, and 3 plot the Entropy Decay Curves for each model–corpus pair. The $x$-axis is logarithmic to highlight behaviour at short context lengths. These curves reveal how each architecture leverages context and where their strategies diverge.
Figure 1: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Alice’s Adventures in Wonderland” corpus. The horizontal axis is shown on a logarithmic scale.
Figure 2: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Ulysses” corpus. The horizontal axis is shown on a logarithmic scale.
Figure 3: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Kant’s Critique of Judgement” corpus. The horizontal axis is shown on a logarithmic scale.
  • Large-Scale Model (Llama 3.3). At very short contexts ($k=3$) Llama 3.3 shows the highest initial uncertainty, signalling a wide search over possible continuations. Its curve then drops steeply, reaching the lowest uncertainty once the context exceeds about $k=90$. This rapid shift from broad exploration to sharp focus appears to be a hallmark of its large parameter count.
  • Smaller Models (DeepSeek-R1 and Qwen 2.5). The two smaller models display a gentler decline. On challenging texts such as Ulysses and Kant’s Critique of Judgement, their curves remain relatively flat and end above $u_{k}=0.35$. This suggests a limited ability to exploit long-range context compared with the larger model.
Corpus-Dependent Profiles and Text Complexity
The uncertainty index $u_{k}$ varies strongly with the corpus. For all models, Alice’s Adventures in Wonderland yields the lowest $u_{k}$ values, indicating that this text is highly predictable. In contrast, Ulysses and Kant’s Critique produce higher $u_{k}$ values, reflecting their greater difficulty for next-token prediction. These differences show that the Entropy Decay Curve (EDC) captures how model capability interacts with linguistic complexity.
4.3 Detecting Memorisation and Data Contamination
The EDC also flags potential memorisation. When Llama 3.3 is run on the Alice corpus, its $u_{k}$ values fall almost to zero at long contexts, an extremity unlikely to arise from genuine generalisation. This strongly suggests that large portions of Alice appeared in its pre-training data. By contrast, on the Kant corpus the same model keeps noticeably higher $u_{k}$ values, consistent with real predictive processing. Thus the EDC can (i) audit test sets for contamination and (ii) separate models that truly generalise from those that rely mainly on memorised content.
4.4 The Information Gain Span (IGS) as a Summary Metric
To summarize the overall trajectory of an entropy decay curve in a single scalar, we introduce the Information Gain Span (IGS):
$\text{IGS}\;:=\;u_{k_{\text{small}}}\cdot\bigl(1-u_{k_{\text{large}}}\bigr),$
where $k_{\text{small}}$ and $k_{\text{large}}$ denote, respectively, short- and long-range context windows. Intuitively, a larger IGS reflects a desirable profile: high initial uncertainty (large $u_{k_{\text{small}}}$) that decays to strong certainty (small $u_{k_{\text{large}}}$). Table 4 reports IGS scores for all experiments using $k_{\text{small}}=3$ and $k_{\text{large}}=600$.
Table 4: Information Gain Span (IGS) Index for Each Model and Corpus. A higher IGS score indicates a better balance between high initial uncertainty and low final uncertainty.
Model Corpus
Alice Ulysses Kant
Llama 3.3 70.6B 0.6764 0.6523 0.5836
DeepSeek-R1 8.19B 0.4339 0.3161 0.3492
Qwen2.5 7.62B 0.5176 0.3225 0.3544
The IGS scores quantitatively confirm our qualitative observations. The Llama 3.3 model achieves the highest IGS score on the Alice and Ulysses corpora, reflecting its superior ability to transition from divergent to convergent processing. All models score lower on the more complex texts, quantitatively demonstrating the increased difficulty.
5 Discussion
The Entropy Decay Curves (EDCs) do more than rank models; they provide a direct, quantitative view of how large language models process information. Because each model–corpus pair yields a distinct curve, the results illuminate properties of both the models and the texts. We highlight two main implications.
5.1 Predictability Profiles of Model–Text Interaction
An EDC serves as a predictability profile: it shows how quickly a model’s uncertainty falls as it reads a text. For every model tested, Ulysses and Kant’s Critique of Judgement produce higher curves than Alice’s Adventures in Wonderland. The first two works therefore pose a longer-lasting predictive challenge.
Unlike traditional readability indices Kincaid et al. (1975), this profile is data-driven. The model itself reveals a text’s effective predictability through its evolving output distribution, without relying on hand-crafted complexity rules.
5.2 Entropy Collapse as an Anomaly Signal
The same method flags unusual behaviour. On the Alice corpus, Llama 3.3 pushes the uncertainty index $u_{k}$ almost to zero at long contexts—a phenomenon we call entropy collapse. Such extreme certainty is unlikely without exposure to the exact text and is therefore strong evidence of memorisation from training-data contamination Carlini et al. (2022); Achiam et al. (2023). Because Alice’s Adventures in Wonderland is a public-domain classic, its presence in pre-training data is plausible.
The pattern differs on the Kant corpus, where Llama 3.3 maintains much higher uncertainty. Thus an entropy collapse on a benchmark text is a reliable warning sign: the model may be recalling rather than generalising, and the training data merit closer inspection.
6 Theoretical Properties of the Entropy Metrics
This section provides a theoretical basis for the empirically observed patterns of $h_{k}$ and $H_{k}$. We move from sample averages to expectations taken with respect to the model’s full probability distribution $p$.
6.1 From Sample Estimates to Expected Values
Definition 1 (Theoretical Entropy Metrics).
The theoretical average conditional entropy is
$h_{k}^{\text{theory}}\;:=\;\mathbb{E}_{X_{1:k}\sim p}\!\bigl[H\!\bigl(p(Y_{k+1}\mid X_{1:k})\bigr)\bigr].$
The theoretical entropy of the average distribution is
$H_{k}^{\text{theory}}\;:=\;H\!\Bigl(\,\mathbb{E}_{X_{1:k}\sim p}\!\bigl[p(Y_{k+1}\mid X_{1:k})\bigr]\Bigr).$
Because the inner expectation is the marginal distribution of $Y_{k+1}$, we have
$H_{k}^{\text{theory}}\;=\;H\!\bigl(p(Y_{k+1})\bigr).$
The empirical $h_{k}$ and $H_{k}$ are Monte Carlo estimates of these theoretical quantities. For the rest of this section we drop the superscript “theory” for clarity.
6.2 Monotonicity of the Average Conditional Entropy ($h_{k}$)
Empirical evidence shows that the average conditional entropy $h_{k}$ does not increase as the context length $k$ grows. A formal proof is subtler, because language is non-stationary and a direct appeal to the inequality $H(Y\mid X,A)\leq H(Y\mid X)$ would compare different prediction targets ($Y_{k+1}$ versus $Y_{k+2}$). To bridge this gap, we first state a sufficient, non-trivial condition that well-trained language models are expected to satisfy. It captures the idea that a model’s confidence, given a fixed context, should not improve when predicting events further in the future.
Proposition 1 (Monotonicity of $h_{k}$ under Predictive Attenuation).
Let $p$ be an autoregressive model. Assume the Predictive Attenuation Condition: for every $k\geq 0$,
$\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr]\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+1}\mid X_{1:k})\bigr].$ (2)
Then $h_{k+1}\leq h_{k}$.
Proof.
Step 1. For any fixed context $X_{1:k+1}$,
$H(Y_{k+2}\mid X_{1:k},X_{k+1})\;\leq\;H(Y_{k+2}\mid X_{1:k}).$
Taking expectations over $X_{1:k+1}$ yields
$h_{k+1}\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr].$ (3)
Step 2. Applying (2),
$\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr]\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+1}\mid X_{1:k})\bigr]\;=\;h_{k}.$ (4)
Combining (3) and (4) gives $h_{k+1}\leq h_{k}$. ∎
Why the Condition Is Reasonable.
Equation (2) states that, on average, a fixed context tells the model less about a token two steps ahead than about the next token. Because the unobserved variable $Y_{k+1}$ introduces additional uncertainty, this behaviour is expected in well-trained language models.
6.3 Monotonicity of the Marginal Entropy ($H_{k}$)
Models trained on natural language typically show a steady decline in $H_{k}$, but this is not an information-theoretic necessity. The trend reflects the non-stationary character of language.
Remark 1 (Why $H_{k}$ Depends on $k$).
We have $H_{k}=H(p(Y_{k+1}))$. If $p(Y_{j})$ were the same for every position $j$, then $H_{k}$ would not change with $k$; a process with position-independent statistics is stationary. Natural language is not stationary: word probabilities differ by position (e.g., first words differ from tenth words). Language models learn this and therefore satisfy $p(Y_{k+1})\neq p(Y_{j+1})$ in general. Hence $H_{k}$ must change with $k$. The observed decline of $H_{k}$ means that as context grows the set of plausible next tokens shrinks, so the marginal distribution becomes more predictable and its entropy falls.
A Bayesian View of the Entropy Metrics.
We can read the three entropy-based quantities as follows:
  • Prior uncertainty: $H_{k}=H(Y_{k+1})$ is the uncertainty before seeing a specific context $X_{1:k}$.
  • Expected posterior uncertainty: $h_{k}=\mathbb{E}[H(Y_{k+1}\mid X_{1:k})]$ is the expected uncertainty after observing the context.
  • Expected information gain: $I_{k}:=I(X_{1:k};Y_{k+1})$ is the expected drop in uncertainty, i.e. the mutual information.
They satisfy the identity
$H_{k}\;=\;h_{k}+I_{k}.$
Proposition 2 (Monotonicity of $H_{k}$ under Predictive Convergence).
If an autoregressive model obeys the Axiom of Predictive Convergence
$h_{k+1}+I_{k+1}\;\leq\;h_{k}+I_{k}\qquad\text{for all }k\geq 0,$
then $H_{k+1}\leq H_{k}$.
Proof.
The identity $H_{j}=h_{j}+I_{j}$ makes the axiom equivalent to $H_{k+1}\leq H_{k}$. ∎
Interpretation.
The axiom says that the model’s total predictive uncertainty cannot grow as it sees more data. Rewriting it as $I_{k+1}-I_{k}\leq h_{k}-h_{k+1}$ yields the Axiom of Diminishing Information Efficiency: the extra mutual information from one more token never exceeds the drop in expected posterior uncertainty.
6.4 Relationship Between the Decay Rates of $h_{k}$ and $H_{k}$
Define single-step drops $\Delta h_{k}:=h_{k}-h_{k+1}\geq 0$ and $\Delta H_{k}:=H_{k}-H_{k+1}\geq 0$.
Remark 2 (Two Regimes of Uncertainty Reduction).
The ratio of $\Delta h_{k}$ to $\Delta H_{k}$ shows what kind of information the new token provides:
  1. 1.
    Local-prediction dominance ($\Delta h_{k}>\Delta H_{k}$). The token greatly helps predict the very next symbol, so $h_{k}$ drops sharply, but it does little to reshape the long-term distribution, so $H_{k}$ falls less.
  2. 2.
    Global-structure dominance ($\Delta H_{k}>\Delta h_{k}$). The token clarifies the topic or style, strongly narrowing the whole future space. The Global Structure Dominance Condition is $\Delta H_{k}\geq\Delta h_{k}$.
Which regime applies can vary with model, corpus, and $k$. Comparing these decay rates is therefore a useful diagnostic.
6.5 A Symmetric View via Bayesian Decomposition
Proposition 3 (Bayesian Decomposition of Marginal Entropy).
For every $k\geq 1$,
$H_{k}-h_{k}\;=\;H(X_{1:k})-H(X_{1:k}\mid Y_{k+1}).$
Proof.
Mutual information symmetry gives
$H(A)-H(A\mid B)\;=\;H(B)-H(B\mid A).$
Let $A:=Y_{k+1}$ and $B:=X_{1:k}$. Then $H_{k}=H(Y_{k+1})$ and $h_{k}=H(Y_{k+1}\mid X_{1:k})$, yielding the stated identity. ∎
Interpretation.
$H_{k}-h_{k}=I(X_{1:k};Y_{k+1})$ measures forward predictive power, while the right-hand side measures how much the next token tells us about the past. They are equal because mutual information is symmetric.
6.6 Theoretical Analysis of the Information-Gain Span
Let a short and long context window be $k_{\mathrm{s}}\ll k_{\mathrm{l}}$. Recall the Information-Gain Span
$\operatorname{IGS}(k_{\mathrm{s}},k_{\mathrm{l}}):=u_{k_{\mathrm{s}}}\bigl(1-u_{k_{\mathrm{l}}}\bigr),\hskip 17.00024ptu_{k}:=h_{k}/H_{k}.$
Theorem 1 (IGS Peak and Markov Order).
Assume the text source is an ergodic Markov chain of exact order $m$ and the model predicts perfectly. Assume further that $H_{k}$ is non-increasing. Then $\operatorname{IGS}(k_{\mathrm{s}},k_{\mathrm{l}})$ is maximised when
$k_{\mathrm{s}}\leq m<k_{\mathrm{l}}.$
Proof.
Step 1: Behaviour of $u_{k}$. For $k<m$, $h_{k}$ falls strictly and $H_{k}$ does not rise, so $u_{k}$ falls. For $k\geq m$, $h_{k}$ is constant ($h_{m}$) while $H_{k}$ still falls, so $u_{k}$ rises or stays flat. Thus $u_{k}$ has a single minimum at $k=m$.
Step 2: The factor $(1-u_{k_{\mathrm{l}}})$. This term is largest when $u_{k_{\mathrm{l}}}$ is smallest, i.e. when $k_{\mathrm{l}}$ is just beyond $m$ (say $m+1$).
Step 3: The factor $u_{k_{\mathrm{s}}}$. Because $u_{k}$ falls on $k<m$, $u_{k_{\mathrm{s}}}$ is largest when $k_{\mathrm{s}}$ is as small as allowed, subject to $k_{\mathrm{s}}\leq m$.
The product is therefore maximised when $k_{\mathrm{s}}$ is small ($\leq m$) and $k_{\mathrm{l}}$ is just above $m$. ∎
Interpretation.
The theorem shows that if a model truly learns a Markov order $m$, then the IGS curve peaks where the short window lies inside that horizon and the long window lies just outside it. By scanning $k_{\mathrm{s}}$ and $k_{\mathrm{l}}$, we can locate the peak and treat its position as an estimate of the model’s effective memory.
7 Conclusion
We introduced a framework for characterising the information–processing behaviour of large language models (LLMs) via their Entropy Decay Curves (EDCs), obtained by plotting the length-conditional uncertainty index $u_{k}$ against context length $k$. This approach yields dynamic cognitive profiles that go beyond conventional single-value metrics.
Empirically, these profiles vary markedly with (i) model architecture, (ii) parameter scale, and (iii) the intrinsic predictability of the input text. The distinct decay trajectories traced by Llama 3.3, DeepSeek-R1, and Qwen 2.5 reveal different strategies for exploiting contextual information. We further defined the Information Gain Span (IGS), a concise statistic that summarises the overall shape of an EDC and enables direct quantitative comparison between models.
The method also functions as a diagnostic tool. The “entropy collapse” observed for Llama 3.3 on the Alice corpus provides strong quantitative evidence consistent with memorisation arising from pre-training data contamination.
The metrics rest on solid theory. First, we proved that the average conditional entropy $h_{k}$ decreases monotonically under a Predictive Attenuation Condition. Second, we showed that the IGS peaks precisely when additional context no longer improves predictive accuracy.
Future work will apply this methodology to a wider range of models and specialised corpora (e.g., code, medical, legal) to chart how cognitive profiles vary with training domain, pushing toward a deeper, mechanistic understanding of artificial intelligence.
Acknowledgements
This study relies exclusively on publicly released large language models. We thank the developers and the wider open-source community for providing these resources.
All experiments used three models: a 70.6B-parameter member of the Llama 3 family, DeepSeek-R1, and Qwen 2.5. In accordance with the licence of the first model, we state: Built with Llama.
Licensing.
  • Llama 3.3: distributed under the Llama 3.3 Community License. Copyright © Meta Platforms, Inc. All rights reserved.
  • DeepSeek-R1: released under the MIT Licence.
  • Qwen 2.5: released under the Apache Licence, Version 2.0.
Our use of each model complies fully with its licence and acceptable-use policy.
This work was partially supported by the KIST Institutional Program.